how many molecules are there in 19.99 g of H2O2?


1
Expert's answer
2021-01-12T07:04:06-0500

Solution:

Moles of H2O2 = Mass of H2O2 / Molar mass of H2O2

The molar mass of H2O2 is 34.0147 g mol-1.

Moles of H2O2 = 19.99 g / 34.0147 g mol-1 = 0.5877 mol


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of H2O2 molecules = 0.5877 mol × (6.022×1023 molecules / 1 mol) = 3.539×1023 molecules

Number of H2O2 molecules = 3.54×1023 molecules


Answer: 3.54×1023 molecules are there in 19.99 g of H2O2.